# BadPart: Unified Black-box Adversarial Patch Attacks against Pixel-wise Regression Tasks

### v1

@@ 6 Conclusion
7 Broader Impact Statements
The unified adversarial patch attack we proposed against pixel-wise regression models aims to disclose the vulnerabilities in such models under query-based black-box attacks. Our work highlights potential security risks in applications that rely on those models, such as autonomous driving, virtual reality and video compositions. We hope to draw the attention of the related developers, and motivate the machine learning (ML) community to create more robust models or defense mechanisms against these types of attacks. This study around the robustness of models is aligned with many prior works/attacks in the ML community, and aims to advance the field of ML. Nevertheless, it is also worth mentioning that our technique can be used for benign purposes, like privacy protection, as we discussed in §5.5.
@@ Appendix A Experimental Details
Attack Settings. Adversarial patches are generated utilizing a single GPU (Nvidia RTX A6000) equipped with a memory capacity of 48G, in conjunction with an Intel Xeon Silver 4214R CPU. The resolution of input scenes from the KITTI dataset is resized to $384\times 1280$ for Planedepth and $320\times 1024$ for other models. We establish the initial square area as 2.5% of the patch area, and the pre-defined square size schedule (Algorithm 2 line 4) is set at $100,500,1500,3000,5000,10000$ for a maximum of 10000 iterations. The square area is halved once the iteration index reaches the pre-defined steps. The initial noise bound $\alpha$ (Algorithm 1 line 7) and noise decay factor $\gamma$ (Algorithm 1 line 23) are set to $0.1$ and $0.98$ respectively. The initialization period $K$ (Algorithm 2 line 5) is $100$ iterations. We adopt an Adam optimizer with the learning rate of $0.1$, and set $0.5$ for both $\beta_{1}$ and $\beta_{2}$. The reference white-box attack in Table 1 also employs the same Adam optimizer, while the gradients for the patch region are calculated through back-propagation. Other hyper-parameters are discussed in the ablation studies, in which we use $b=20$, $T_{1}=1$ and $T_{2}=1$ as the default settings.
@@ Appendix B Additional Ablation Studies
Models GA HB S-RS Ours GA HB S-RS Ours GA HB S-RS Ours
@@ Appendix C More Qualitative and Quantitative Results
Appendix D Limitations and Future Work
@@
In this work, we have explored the black-box adversarial patch attack against pixel-wise regression models, which reveals the potential vulnerabilities in such models and their expanded applications. We have addressed the domain-specific challenge of high-resolution patch optimization, and our proposed method has shown an attack efficacy that surpasses that of various established benchmarks. It also appears to be robust against state-of-the-art black-box defense mechanisms. However, it is important to acknowledge potential limitations of our study. Our focus was predominantly on digital-space attacks, wherein the perturbed pixels are utilized directly as input for the model. This approach aligns with the conventional methods employed in prior black-box attacks, as referenced in works such as (Andriushchenko et al., 2020; Croce et al., 2022; Tao et al., 2023), and is deemed practical within the threat model that encompasses attacks on online services. Nevertheless, the implications of physical-world attacks are arguably more profound, particularly in the context of autonomous driving. Physical-world attacks, such as those described in (Cheng et al., 2022), necessitate a more nuanced consideration of environmental factors, including but not limited to viewing angles, distances, lighting conditions, and printing qualities. The question of how to effectively execute query-based black-box patch attacks in a physical setting remains unresolved. It is this intriguing challenge that we aim to address in our future research endeavors. An additional constraint pertains to the dimensions of the adversarial patch. Should individuals seek to employ such patches for the purpose of privacy preservation, as discussed in §5.5, the current manifestation of the patch on portrait images may be conspicuously apparent. The generation of stealthy patches for pixel-wise regression tasks within a black-box context is an unresolved issue, and we earmark this as a topic for future investigation.

### v3

@@ 6 Conclusion
Impact Statement
The unified adversarial patch attack that we proposed against pixel-wise regression models aims to disclose the vulnerabilities in such models under query-based black-box attacks. Our work highlights potential security risks in applications that rely on those models, such as autonomous driving, virtual reality, and video compositions. We hope to draw the attention of the related developers, and motivate the machine learning (ML) community to create more robust models or defense mechanisms against these types of attacks. This study around the robustness of models is aligned with many prior works/attacks in the ML community, and aims to advance the field of ML. Nevertheless, it is also worth mentioning that our technique can be used for benign purposes, such as protecting privacy, as we discussed in §5.5.
@@ Appendix A Experimental Details
Attack Settings. Adversarial patches are generated utilizing a single GPU (Nvidia RTX A6000) equipped with a memory capacity of 48G, in conjunction with an Intel Xeon Silver 4214R CPU. The resolution of input scenes from the KITTI dataset is resized to $384\times 1280$ for Planedepth and $320\times 1024$ for other models. We establish the initial square area as 2.5% of the patch area, and the pre-defined square size schedule (Algorithm 2 line 4) is set at $100,500,1500,3000,5000,10000$ for a maximum of 10000 iterations. The square area is halved once the iteration index reaches the pre-defined steps. The initial noise bound $\alpha$ (Algorithm 1 line 7) and noise decay factor $\gamma$ (Algorithm 1 line 23) are set to $0.1$ and $0.98$ respectively. The initialization period $K$ (Algorithm 2 line 5) is $100$ iterations. We adopt an Adam optimizer with the learning rate of $0.1$, and set $0.5$ for both $\beta_{1}$ and $\beta_{2}$. The reference white-box attack in Table 1 also employs the same Adam optimizer, while the gradients for the patch region are calculated through back-propagation. Other hyper-parameters are discussed in the ablation studies, in which we use $b=20$, $T_{1}=1$ and $T_{2}=1$ as the default settings. We discuss the transferability of our approach to another dataset in Appendix E.
@@ Appendix B Additional Ablation Studies
Models GA HB P-RS Ours GA HB P-RS Ours GA HB P-RS Ours
@@ Appendix C More Qualitative and Quantitative Results
Appendix D Discussion on More Attacks and Defenses
Attacks. In our main experimetns, we have endeavored to compare our method with the SOTA score-based (i.e., soft-label) attack, specifically Patch-RS (Croce et al., 2022), and decision-based (i.e., hard-label) attack, namely HardBeat (Tao et al., 2023). Given the new challenges encountered in black-box patch attacks against pixel-wise regression models, we had to adapt these SOTA attacks which were originally designed for classification models, to ensure an equitable comparison. It is important to note, however, that not every method developed for classification models is amenable to adaptation for our scenario. This is, in part, due to the absence of class labels in our context, which are integral to some methods. For instance, PatchAttack (Yang et al., 2020) leverages pre-generated texture images for each class in ImageNet, which are imbued with distinctive class features capable of triggering the corresponding label, generated using a white-box surrogate model. Similarly, DevoPatch (Chen et al., 2023) utilizes an image from the target class (e.g., an elephant) as foundational prior knowledge to initiate the patch on a source image (e.g., a clock). The reliance on class labels in these methods renders them incompatible with our investigation of purely query-based black-box attacks in the realm of pixel-wise regression tasks.
HPA (Fawzi & Frossard, 2016) and Adv-watermark (Jia et al., 2020) are also designed for classification tasks. However, HPA (Fawzi & Frossard, 2016) primarily utilizes the classifier’s output logits, which we can adapt to our context by using our regression outputs. Meanwhile, Adv-watermark (Jia et al., 2020) employs an evolutionary algorithm to optimize the positioning and transparency of watermarks that serve as adversarial patches. To evaluate the attack performance of HPA and Adv-watermark on pixel-wise regression tasks, we conducted additional experiments, and results are shown in Table 9. As shown, the errors caused by those attacks on both MDE and OFE tasks are limited, and additional queries do not significantly enhance their attack performance, suggesting low efficiency. The performance of Adv-watermark remains nearly static after initialization. In contrast, our approach proves to be substantially more effective and efficient.
Table 9: Attack performance comparison with additional attack baselines (single-image).
Monodepth2 FlowNet2
Query HPA Adv-watermark BadPart (ours) HPA Adv-watermark BadPart (ours)
10K 1.376 2.178 9.060 3.024 4.551 13.141
30K 1.386 2.178 24.235 3.335 4.551 41.562
50K 1.409 2.178 50.953 3.542 4.553 53.663
80K 1.503 2.178 70.598 3.677 4.553 60.793
100K 1.598 2.178 76.890 3.828 4.554 67.954
It is also pertinent to mention that the above attacks are all single-image attacks, in which each image requires a unique patch pattern optimized specifically. However, we focus on the universal attack, in which the patch generation is a one-time effort and the generated patch can be applied to arbitrary unseen images and attack universally. This fundamental difference in problem setting also elucidates the impracticality of applying the aforementioned methods to our scenario.
Defenses. We have also further analyzed the mainstream defense methods against patch attacks and have ported the applicable defense algorithms to our scenario to test the effectiveness of our attack methods. The Local Gradient Smoothing (LGS) algorithm (Naseer et al., 2019) implements security defense through identifying the high-frequency patch areas and utilizing local gradient smoothing to degrade the patch’s effectiveness. However, in our experiments, this method results in a high rate of false positive anomaly identification and smoothing on benign areas. Although the false positive smoothing may not affect the classification tasks, it severely affected the pixel-wise regression outputs of our subject models in areas that are incorrectly smoothed. Detailed results can be found in Table 10. The algorithm’s parameters are set to the optimal configuration as per the paper. As shown, the defense affects, on average, 25.245% of the total area in randomly selected benign images, leading to an average relative prediction error of 23.344% across various pixel-wise regression models. This renders the defense impractical for real-world application.
Table 10: Model performance degradation on benign images caused by LGS (Naseer et al., 2019).
Models Affected Portion (%) Relative Error (%)
Monodepth2 27.335 16.698
DepthHints 25.944 17.010
PlaneDepth 31.781 15.738
SQLDepth 35.982 18.401
FlowNetC 19.037 29.740
FlowNet2 19.695 27.507
PWC-Net 16.939 38.316
Average 25.245 23.344
Moreover, many defense methods against patch attacks for classification tasks are not applicable in our scenario of pixel-wise regression tasks. DW (Hayes, 2018) treats defense as a process similar to watermark removal targeted at patches. Although the guided back-propagation method mentioned in this article for constructing saliency maps of images might help in effectively identifying patch areas, the algorithm relies on the use of predicted labels, which do not exist in our scenario. RS (Levine & Feizi, 2020) defends against patch attacks on classification models effectively using statistical predictions on a small number of image pixels. However, this practice is based on the fact that classification models could still predict the right label using only a fraction of the image. In contrast, pixel-wise regression models cannot make accurate predictions on partial areas; they need to process the entire scene, which makes this defense strategy not applicable. Based on our analysis and experiments on defense algorithms against patch attacks, there currently exists no defense method against our method that can balance good defense effectiveness with small impact on the performance of benign samples. The protection of pixel-wise regression models against attacks remains a significant challenge.
Appendix E Transferability of the Dataset
The pre-training process in Algorithm 1 is different from model training and we do not require access to the model’s training images. Our attack is the universal patch attack instead of single-image attack. It requires a one-time patch generation process (Algorithm 1), referred to as the “pre-training stage”, then the generated patch can be applied to arbitrary unseen images to attack the target model. As stated in the Attack Settings of Section 5.1, the attack performance in our experiments is evaluated using an unseen test set (also customizable), and Figure 11 shows the qualitative results, demonstrating the universal effectiveness of our patches across various scenes. Therefore, although we call the patch generation process in Algorithm 1 as a “pre-training stage”, the training and validation images used in this stage are customizable. We have also conducted additional experiments in which we use images from another public dataset, named nuScenes, as our training set for patch generation, to attack models trained on the KITTI dataset. Results are shown in Table 11. As shown, patches generated using the nuScenes dataset still achieve a good attack performance, on par with those generated using the KITTI dataset. This validates that access to the training images of the model is not a prerequisite.
Table 11: Attack performance of patches generated using different dataset.
Model Monodepth2 FlowNetC
[…]
50K 28.568 19.439 149.397 165.549
100K 42.541 28.397 208.646 270.407
200K 65.092 45.208 313.976 492.197
300K 78.380 66.210 503.938 695.503
Since our attack is dataset-independent and does not require access to the model’s training set, as long as the subject model works on the input image, the generated patch can be effective. We have conducted additional experiments to validate the transferability of the patch generated on KITTI to nuScenes. We sample 20 images randomly from the nuScenes dataset as the test set, and report the attack performance for different models using patches generated with various numbers of queries. The result can be found in Table 12. As shown, the attack performance on images from the KITTI dataset is similar to the attack performance on the nuScenes dataset, which validates that our attack is dataset-agnostic.
Table 12: Attack performance of transferring the patch generated from KITTI to nuScenes dataset.
Model Monodepth2 DepthHints FlowNetC
[…]
50K 20.481 19.439 8.285 6.803 164.573 165.549
200K 48.919 45.208 36.136 30.86 495.177 492.197
400K 80.714 81.097 44.276 42.832 753.558 742.817
800K 85.945 89.321 66.031 60.678 1149.475 1132.284
Appendix F Limitations and Future Work
@@
In this work, we have explored the black-box adversarial patch attack against pixel-wise regression models, which reveals the potential vulnerabilities in such models and their expanded applications. We have addressed the domain-specific challenge of high-resolution patch optimization, and our proposed method has shown an attack efficacy that surpasses that of various established benchmarks. It also appears to be robust against state-of-the-art black-box defense mechanisms. However, it is important to acknowledge potential limitations of our study. Our focus was predominantly on digital-space attacks, wherein the perturbed pixels are utilized directly as input for the model. This approach aligns with the conventional methods employed in prior black-box attacks, as referenced in works such as (Andriushchenko et al., 2020; Croce et al., 2022; Tao et al., 2023), and is deemed practical within the threat model that encompasses attacks on online services. Nevertheless, the implications of physical-world attacks are arguably more profound, particularly in the context of autonomous driving. Physical-world attacks, such as those described in (Cheng et al., 2022; Choi et al., 2022), necessitate a more nuanced consideration of environmental factors, including but not limited to viewing angles, distances, lighting conditions, and printing qualities. The question of how to effectively execute query-based black-box patch attacks in a physical setting remains unresolved. It is this intriguing challenge that we aim to address in our future research endeavors. An additional constraint pertains to the dimensions of the adversarial patch. Should individuals seek to employ such patches for the purpose of privacy preservation, as discussed in §5.5, the current manifestation of the patch on portrait images may be conspicuously apparent. The generation of stealthy patches for pixel-wise regression tasks within a black-box context is an unresolved issue, and we earmark this as a topic for future investigation.